OpenStudy (anonymous):

evaluate: integral from 3 to infinity: 1/(x-2)^(3/2) dx

OpenStudy (anonymous):

\[I=\int_3^\infty\frac{1}{(x-2)^{3/2}}dx\\ u=x-2\implies du=dx\\ \text{when}\;x=3,u=1\\ \text{when}\;x=\infty,u=\infty\\ I=\int_1^\infty\frac{1}{u^{3/2}}du\\ I=\int_1^\infty u^{-3/2}du\\ I=\left[\frac{u^{-1/2}}{-1/2}\right]_1^\infty=-2\left[{1\over\sqrt{u}}\right]_1^\infty\\ I=-2\left({1\over\infty}-1\right)\\ \boxed{I=2} \]